OpenStudy (anonymous):

Given A=3i+4j and B=6u+8j. which of the following statement is incorrect ? a) AXB=20 b)|A|/|b|=1/2 c)|A|=5 c)A.B=48Given A=3i+4j and B=6u+8j. which of the following statement is incorrect ? a) AXB=20 b)|A|/|b|=1/2 c)|A|=5 c)A.B=48@Mathematics

myininaya (myininaya):

\[|A|=\sqrt{3^2+4^2}=\sqrt{9+16}=\sqrt{25}=5\] \[|B|=\sqrt{6^2+8^2}=\sqrt{36+64}=\sqrt{100}=10\] 5/10=1/2 so choice b and c is good \[A \cdot B=<3,4> \cdot <6,8>=3(6)+4(8)=18+32=50 \neq 48\] the second c choice is a bad one ( iassumed you meant B=6i+8j right?)

myininaya (myininaya):

no reason to check the first

myininaya (myininaya):

but choice a is also wrong

myininaya (myininaya):

you are suppose to get another vector